# Positive weak solutions of a double-phase variable exponent problem with a fractional-Hardy-type singular potential and superlinear nonlinearity

### v2

@@ 2 Mathematical Background and Preliminaries
$C_{+}\left(\overline{\Omega}\right)=\left\{h\in C\left(\overline{\Omega}\right),h(x)>1\text{ for all\ }x\in\overline{\Omega}\right\}.$
@@
For any $h\in C_{+}\left(\overline{\Omega}\right)$, we define the variable exponent Lebesgue space by
$L^{h(x)}(\Omega)=\left\{u\mid u:\Omega\rightarrow\mathbb{R}\text{ is measurable},\int_{\Omega}|u(x)|^{h(x)}dx<\infty\right\}.$
Then, $L^{h(x)}(\Omega)$ endowed with the norm
$|u|_{h(x)}=\inf\left\{\lambda>0:\int_{\Omega}\left|\frac{u(x)}{\lambda}\right|^{h(x)}dx\leq 1\right\},$
becomes a Banach space.
Proposition 2.1 (Hölder Inequality).
@@
for all $u\in W^{1,h(x)}(\Omega)$.
@@
If $1<h^{-}\leq h^{+}<\infty$, then the spaces $L^{h(x)}(\Omega)$ and $W^{1,h(x)}(\Omega)$ are separable and reflexive Banach spaces.
The space $W_{0}^{1,h(x)}(\Omega)$ is defined as
$W_{0}^{1,h(x)}(\Omega):=\bigcap_{C_{0}^{\infty}(\Omega)\subset\mathcal{K}}\mathcal{K}\quad\textit{ for closed subsets }\mathcal{K}\subset W^{1,h(x)}(\Omega).$
Therefore, $W_{0}^{1,h(x)}(\Omega)$ is also a separable and reflexive Banach space. Furthermore, since Poincaré inequality holds on $W_{0}^{1,h(x)}(\Omega)$, we have
$|\nabla u|_{h(x)}\leq\|u\|_{1,h(x)}\leq(C+1)|\nabla u|_{h(x)},\quad\forall u\in W_{0}^{1,h(x)}(\Omega),\quad 0<C\in\mathbb{R},$
which implies $\|u\|_{1,h(x)}$ and $|\nabla u|_{h(x)}$ are equivalent norms on $W_{0}^{1,h(x)}(\Omega)$. Therefore, for any $u\in W_{0}^{1,h(x)}(\Omega)$ we can define an equivalent norm $\|u\|$ such that
$\|u\|=|\nabla u|_{h(x)}.$
@@
Let $r\in C(\overline{\Omega})$. If $1\leq r(x)<h^{\ast}(x)$ for all $x\in\overline{\Omega}$, then the embeddings $W^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ and $W_{0}^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ are compact and continuous, where
@@
    $p,q\in C(\overline{\Omega})$, $1<p(x),q(x)<N$ with $p(x)<q(x)<p^{*}(x)$ for all $x\in\overline{\Omega}$.
@@
    $\mu\in L^{\infty}(\Omega)$ such that $\mu(\cdot)\geq 0$.
@@
Analogous to Proposition 2.2, there are similar relationship between the modular $\rho_{\mathcal{H}}(\cdot)$ and the norm $\|\cdot\|_{\mathcal{H}}$, see [17, Proposition 2.13] for a detailed proof. Crespo-Blanco
@@
where $\|\nabla u\|_{\mathcal{H}}=\|\,|\nabla u|\,\|_{\mathcal{H}}$, and $|\nabla u|$ is the $n$-dimensional Euclidean norm of the gradient of $u$.
The space $W_{0}^{1,\mathcal{H}}(\Omega)$ is defined (in a similar way as with $W_{0}^{1,h(x)}(\Omega)$) as $W_{0}^{1,\mathcal{H}}(\Omega):=\overline{C_{0}^{\infty}(\Omega)}^{\|\cdot\|_{1,\mathcal{H}}}$. Notice that $L^{\mathcal{H}}(\Omega),W^{1,\mathcal{H}}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)$ are uniformly convex and reflexive Banach spaces. Moreover we have the following embeddings [17, Proposition 2.16].
@@
Let $(H)$ be satisfied. Then the following embeddings hold:
@@
    $L^{\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega),W^{1,\mathcal{H}}(\Omega)\hookrightarrow W^{1,h(x)}(\Omega)$, $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow W_{0}^{1,h(x)}(\Omega)$ are continuous for all $h\in C(\overline{\Omega})$ with $1\leq h(x)\leq p(x)$ for all $x\in\overline{\Omega}$.
@@
    $W^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ are compact for all $h\in C(\overline{\Omega})$ with $1\leq h(x)<p^{*}(x)$ for all $x\in\overline{\Omega}$.
As a conclusion of Proposition 2.7, we have the continuous embedding $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$, and $C_{\mathcal{H}}$ denotes the best constant such that
$\|u\|_{h(x)}\leq C_{\mathcal{H}}\|u\|_{1,\mathcal{H},0}.$
Moreover, by [17, Proposition 2.18], $W_{0}^{1,\mathcal{H}}(\Omega)$ is compactly embedded in $L^{\mathcal{H}}(\Omega)$. Thus, the space $W_{0}^{1,\mathcal{H}}(\Omega)$ can be equipped with the equivalent norm
$\|u\|_{1,\mathcal{H},0}=\|\nabla u\|_{\mathcal{H}}.$
@@
is of class $C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$, and its derivative $\varrho^{\prime}_{\mathcal{H}}$ satisfies the ($S_{+}$)-property with the derivative
@@
Assume that $(H_{1})$ holds. If, additionally, $\alpha p^{-}\neq 1$ and $\alpha q^{+}\neq 1$, then the following hold:
$(i)$ There exists a positive constant $C_{N}(p^{-},q^{+},\alpha)$ such that the inequality
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-},q^{+},\alpha)\|u\|_{1,\mathcal{H},0}^{\kappa},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega),$ (1)
holds, where $C_{N}(p^{-},q^{+},\alpha):=\frac{(1+\|\mu\|_{\infty})}{p^{-}}\max\{C_{N}(q^{+}),C_{N}(p^{-})\}$ with $C_{N}(q^{+})=\hat{c}_{1}\left(\frac{q^{+}}{N-\alpha q^{+}}\right)^{q^{+}}$, $C_{N}(p^{-})=\hat{c}_{2}\left(\frac{p^{-}}{N-\alpha p^{-}}\right)^{p^{-}}$ where $\hat{c}_{1}>0,\hat{c}_{2}>0$ are the embedding constants not depending on $u$; and $\kappa=p^{-}$ if $\|u\|_{1,\mathcal{H},0}<1$, $\kappa=q^{+}$ if $\|u\|_{1,\mathcal{H},0}\geq 1$.
$(ii)$ There exists $0<M\in\mathbb{R}$ such that the inequality
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx>\frac{1}{M^{\alpha\tau}}\|u\|^{\kappa}_{\mathcal{H}},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega)$ (2)
holds, where $\tau=p^{-}$ if $0<|x|<M<1$, and $\tau=q^{+}$ if $1\leq|x|<M$.
@@
$(i)$ We split $\Omega$ in the following two subsets
$\Omega_{<}:=\{x\in\Omega:|x|<1\},\quad\Omega_{\geq}:=\{x\in\Omega:|x|\geq 1\}.$ (3)
Since $\Omega$ is bounded and connected, we obviously have $\Omega=\Omega_{<}\cup\Omega_{\geq}$.
Case I: $x\in\Omega_{<}$.
In this case, we have $|x|^{\alpha}>|x|$. Thus, we need to consider the following orderings:
  • (a)
    $|u(x)|\geq|x|^{\alpha}>|x|$,
  • (b)
    $|x|^{\alpha}>|u(x)|>|x|$,
  • (c)
    $|x|^{\alpha}>|x|\geq|u(x)|$.
For (a): $|u(x)|\geq|x|^{\alpha}$. Using $(H_{1})$, for all $x\in\Omega_{<}$, $x\neq 0$, we have
$\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\leq\frac{1}{p^{-}}\left(\frac{|u(x)|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u(x)|^{q(x)}}{|x|^{\alpha q(x)}}\right)\leq\frac{\left(1+\|\mu\|_{\infty}\right)}{p^{-}}\frac{|u(x)|^{q^{+}}}{|x|^{\alpha q^{+}}}.$ (4)
Henceforth, we follow the same reasoning as with [3].
We start with the identity (adapted from Theorem 7, [20])
$\frac{|u(x)|^{q^{+}}}{|x|^{\alpha q^{+}}}=\frac{1}{(1-\alpha q^{+})}|u(x)|^{q^{+}}\frac{x}{|x|}\cdot\nabla\left[\frac{1}{|x|^{\alpha q^{+}-1}}\right].$ (5)
Using the divergence theorem gives
$\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx=\frac{1}{(\alpha q^{+}-1)}\int_{\Omega_{<}}\nabla\cdot\left[|u|^{q^{+}}\frac{x}{|x|}\right]\frac{1}{|x|^{\alpha q^{+}-1}}dx+\frac{1}{(1-\alpha q^{+})}\int_{\partial\Omega}|u|^{q^{+}}\frac{x}{|x|^{\alpha q^{+}}}\cdot\textbf{n}dS,$ (6)

### v3

@@ 2 Mathematical Background and Preliminaries
$C_{+}\left(\overline{\Omega}\right)=\left\{h\in C\left(\overline{\Omega}\right):1<h(x)\text{ for all\ }x\in\overline{\Omega}\right\}.$
@@
For a given $h\in C_{+}\left(\overline{\Omega}\right)$, we define the variable exponent Lebesgue space by
$L^{h(x)}(\Omega)=\left\{u\mid u:\Omega\rightarrow\mathbb{R}\text{ is measurable},\int_{\Omega}|u(x)|^{h(x)}dx<\infty\right\},$
equipped with the Luxemburg norm given by
$|u|_{h(x)}=\inf\left\{\lambda>0:\int_{\Omega}\left|\frac{u(x)}{\lambda}\right|^{h(x)}dx\leq 1\right\}.$
With this norm, $(L^{h(x)}(\Omega),|\cdot|_{h(x)})$ becomes a separable and reflexive Banach space.
Proposition 2.1 (Hölder’s Inequality).
@@
[…]
The space $W_{0}^{1,h(x)}(\Omega)$ is defined as the closure of $C_{0}^{\infty}(\Omega)$ in $W^{1,h(x)}(\Omega)$.
@@
If $h\in C_{+}(\overline{\Omega})$ and $h^{+}<\infty$, then the spaces $L^{h(x)}(\Omega)$, $W^{1,h(x)}(\Omega)$, and $W_{0}^{1,h(x)}(\Omega)$ are separable and reflexive Banach spaces.
Furthermore, Poincaré inequality holds in $W_{0}^{1,h(x)}(\Omega)$ [21]; that is, there exists a positive constant $c$ independent of $u$ such that
$|u|_{h(x)}\leq c|\nabla u|_{h(x)},\quad\forall u\in W_{0}^{1,h(x)}(\Omega),$
which implies that $|\nabla u|_{h(x)}$ is an equivalent norm in $W_{0}^{1,h(x)}(\Omega)$. Therefore, on $W_{0}^{1,h(x)}(\Omega)$ we can define an equivalent norm $\|\cdot\|_{h(x)}$ such that
$\|u\|_{h(x)}=|\nabla u|_{h(x)}.$
@@
If $h\in C_{+}(\overline{\Omega})$, $r\in C(\overline{\Omega})$ and $1\leq r(x)<h^{\ast}(x)$ for all $x\in\overline{\Omega}$, then the embeddings $W^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ and $W_{0}^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ are compact and continuous, where
@@
    $p,q\in C_{+}(\overline{\Omega})$, $1<p(x)<N$ and $p(x)<q(x)$ for all $x\in\overline{\Omega}$ with
    $q^{+}<\min\{p^{*}(x),N\}$.
@@
    $\mu\in L^{\infty}(\Omega)$ such that $\mu(\cdot)\geq 0$, and $\mu(\cdot)\not\equiv 0$.
@@
Analogous to Proposition 2.2, there are similar relationship between the modular $\rho_{\mathcal{H}}(\cdot)$ and the norm $\|\cdot\|_{\mathcal{H}}$, see [17, Proposition 2.13] for a detailed proof.
@@
where $\|\nabla u\|_{\mathcal{H}}=\|\,|\nabla u|\,\|_{\mathcal{H}}$.
The space $W_{0}^{1,\mathcal{H}}(\Omega)$ is defined as the closure of $C_{0}^{\infty}(\Omega)$ in $W^{1,\mathcal{H})}(\Omega)$. Note also that $L^{\mathcal{H}}(\Omega),W^{1,\mathcal{H}}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)$ are reflexive Banach spaces [17, Proposition 2.12].
We now present the following embedding relations given in [17, Proposition 2.16].
@@
Assume that $(H_{1})$ and $(H_{2})$ hold. Then the following embeddings hold:
@@
    $L^{\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega),W^{1,\mathcal{H}}(\Omega)\hookrightarrow W^{1,h(x)}(\Omega)$, $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow W_{0}^{1,h(x)}(\Omega)$ are continuous for $h\in C(\overline{\Omega})$ with $1\leq h(x)\leq p(x)$ for all $x\in\overline{\Omega}$.
@@
    $W^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ are compact for $h\in C(\overline{\Omega})$ with $1\leq h(x)<p^{*}(x)$ for all $x\in\overline{\Omega}$.
Proposition 2.8.
[17] Assume that $(H_{1})$ and $(H_{2})$ hold. Then the following hold:
  • $(i)$
    The embedding $W^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{\mathcal{H}}(\Omega)$ is compact;
  • $(ii)$
    There exists a constant $c>0$ independent of $u$ such that
    $\|u\|_{\mathcal{H}}\leq c\|\nabla u\|_{\mathcal{H}},\quad\forall u\in W_{0}^{1,\mathcal{H}}(\Omega).$
As a conclusion of Proposition 2.8, the space $W_{0}^{1,\mathcal{H}}(\Omega)$ can be equipped with an equivalent norm $\|\cdot\|_{1,\mathcal{H},0}$ given by
$\|u\|_{1,\mathcal{H},0}=\|\nabla u\|_{\mathcal{H}},\quad\forall u\in W_{0}^{1,\mathcal{H}}(\Omega).$
@@
is of class $C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$, and its derivative $\varrho^{\prime}_{\mathcal{H}}$ satisfies the ($S_{+}$)-property [35] with the derivative
@@
Assume that $(H_{1}),(H_{2})$ hold. If, additionally, $\alpha p^{-}>1$ and $\alpha q^{+}>1$, then the following hold:
$(i)$ There exist constants $\hat{H},\kappa$ independent of $u$ such that the inequality
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq\hat{H}\|u\|_{1,\mathcal{H},0}^{\kappa}$ (1)
holds for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$.
$(ii)$ There exists a constant $M>1$ such that $|x|<M$ for all $x\in\Omega\setminus\{0\}$. Then,
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx>\frac{1}{M^{\alpha q^{+}}}\|u\|^{{p^{-}\wedge q^{+}}}_{\mathcal{H}},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega),$ (2)
where $p^{-}\wedge q^{+}=\min\{p^{-},q^{+}\}$ determined according to the magnitude of $u$.
@@
$(i)$ First, we obtain a fractional-Hardy-type inequality. To do so, we consider a general constant exponent $s$ with $1<s<N$.
For $u\in C_{0}^{\infty}(\Omega)$, we have
$\nabla\cdot\left(\frac{x}{|x|^{\alpha s}}\right)=\frac{N-\alpha s}{|x|^{\alpha s}}.$ (3)
Hence,
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx=\frac{1}{N-\alpha s}\int_{\Omega}\mu(x)\,|u|^{s}\,\nabla\cdot\left(\frac{x}{|x|^{\alpha s}}\right)dx.$ (4)
Applying the Divergence Theorem along with integrating by parts provides
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx$ $=\frac{1}{N-\alpha s}\int_{\Omega}\mu(x)\,|u|^{s}\,\nabla\cdot\left(\frac{x}{|x|^{\alpha s}}\right)dx$
$=-\frac{1}{N-\alpha s}\int_{\Omega}\mu(x)\,\nabla(|u|^{s})\cdot\frac{x}{|x|^{\alpha s}}dx,$ (5)